# See pages - of preprint.docx.pdf

### v2

@@
Recall that during training an amputated module is fully removed; however, during testing the precise location of amputation is uniformly sampled between the conjunction of the torso and each limb and the sphere of that limb. To quantify the performance of the amputation-agnostic policy compared to the original controller across previously unseen amputation locations, 10 sets of cutpoints for each amputation topology were sampled. Each set of remnant configurations was simulated for five seconds, five independent times under both the amputation-agnostic and original control polices. In Fig. 5H-K, the mean and standard deviation of displacement of each amputation scenario is reported across all constituent trials (H: no damage, I: four modules remaining, J: three modules remaining: K: one module remaining). The trajectories of the first of the five simulation repetitions for each of the sampled post-amputation metamachines under the amputation-agnostic policy (blue to green traces) and the original quadruped controller (pink to cyan traces) are shown in Fig. S3 (two rows per amputated topology). In most cases, the amputation-agnostic policy outperforms the original quadruped walking policy.
The parameters of the testing simulation are set to match the real-world experiment, which are shown in Table S10.
@@
Motor position offset [m] $-0.0025$ 0.0025
@@
Term Symbol Weight
Forward velocity $r_{\text{forward}}$ 0.6
Angular velocity $r_{\text{walk-turn}}$ 0.2
Action rate $r_{\text{action}}$ $-0.1$
Num joint fall $r_{\text{fall}}$ $-0.02$
Vel penalty $r_{\text{motor-vel}}$ $-0.01$
Acc penalty $r_{\text{motor-acc}}$ $-0.000002$
@@
Initial torso joint position [radian] 0.6 $-0.6$ 1 $-1$

### v3

@@
Recall that during training an amputated module is fully removed; however, during testing the precise location of amputation is uniformly sampled between the conjunction of the torso and each limb and the sphere of that limb. To quantify the performance of the amputation-agnostic policy compared to the original controller across previously unseen amputation locations, 10 sets of cutpoints for each amputation topology were sampled. Each set of remnant configurations was simulated for five seconds, five independent times under both the amputation-agnostic and original control polices. In Fig. 5H-K, the mean and standard deviation of displacement of each amputation scenario is reported across all constituent trials (H: no damage, I: four modules remaining, J: three modules remaining: K: one module remaining). The trajectories of the first of the five simulation repetitions for each of the sampled post-amputation robots under the amputation-agnostic policy (blue to green traces) and the original quadruped controller (pink to cyan traces) are shown in Fig. S3 (two rows per amputated topology). In most cases, the amputation-agnostic policy outperforms the original quadruped walking policy. The parameters of the testing simulation are set to match the real-world experiment, which are shown in Table S10.
@@
Figure S7: Overview of the design pipeline. Five hundred thousand example designs were randomly generated. Each design is a different configuration tree of at least three and no more than five modules, and is collision-free in its natural pose. A Variational Autoencoder (VAE) was trained to compress this dataset into a latent multivariate normal distribution, characterized by a mean vector ($\mu$) and standard deviations ($\sigma$). Training of the VAE minimizes two objectives: reconstruction loss, which measures how well decoded designs match the originals, and Kullback-Leibler (KL) divergence loss, which regularizes the latent distribution toward a standard normal prior. Bayesian optimization (BO) was conducted within the resulting latent space. BO builds a probabilistic surrogate model (Gaussian process) of the performance landscape, and uses this to propose which latent vectors to evaluate next, i.e. the latent coordinates that are predicted to be within promising unexplored regions of the latent space. Each latent vector is decoded into a configuration tree, instantiated in a physics simulation, and trained to locomote using deep reinforcement learning (RL). The performance of the design was taken to be the average accumulated reward across the final 10% of training episodes. This performance score is then fed back to BO to update the Gaussian process. In RL, an Actor network generates actions from states, while a Critic network estimates Q-values (expected future rewards) to update the Actor network. Optimized Actor policies were deployed on real robots for zero-shot control.
Figure S8: Overview of amputation-agnostic policy training and testing. In each training scenario, different combinations of limbs were amputated from the quadruped. For each scenario, a scenario-specific “expert” policy was trained to control the reduced morphology using reinforcement learning (RL). The sensorimotor sequences generated by the expert policy were collected for each scenario and pooled across scenarios in order to train a single transformer-based controller, an amputation-agnostic policy that generalizes across amputations. At test time, the amputation-agnostic policy was evaluated on diverse, unseen damage scenarios: novel amputation topologies, novel cut positions, and unexpected motor failures.
Figure S9: The utility of the latent design space. In the main experiment, bayesian optimization (BO) operates within a learned latent compression of configuration space (Fig. S7). To determine if this latent space is more easily searchable than the raw (uncompressed) configuration space, an otherwise equivalent experiment was conducted in which BO operates directly on the configuration trees (their one-hot encoding) instead of within the latent space (see Sect. S4.8 for details). Twelve independent trials of BO were conducted, six in the latent space (blue) and six on the raw configuration trees (red). In each trial, the performance of the best design found so far (i.e. cumulative max performance) was recorded across optimization time. These data are summarized with shaded regions indicating the standard error of the mean. On average, BO in latent space discovered significantly better designs than BO in the raw configuration space ($p<0.01$; t test). The performance of the manually designed quadruped is also shown for reference (gray dashed line). In all 12 trials, BO discovered designs with higher performance than the quadruped.
@@
Motor position offset [m] 0.0025-0.0025 0.0025
@@
Term Symbol Weight Weight (symmetrical configurations)
Forward velocity $r_{\text{forward}}$ 0.8 0.6
Angular velocity $r_{\text{walk-turn}}$ 0.2 0.2
Action rate $r_{\text{action}}$ 0 0.1-0.1
Num joint fall $r_{\text{fall}}$ 0.01-0.01 0.02-0.02
Vel penalty $r_{\text{motor-vel}}$ 0.005-0.005 0.01-0.01
Acc penalty $r_{\text{motor-acc}}$ 0.000001-0.000001 0.000002-0.000002
@@
Initial torso joint position [radian] 0.6 0.6-0.6 1 1-1